OpenStudy (anonymous):

Translate this problem to an equation and solve. The sum of the squares of two consecutive positive even numbers is 52. What is the smaller number?

OpenStudy (anonymous):

@jojo4eva @Jwoodward48 @FlvsGirl @rose21

OpenStudy (anonymous):

Actually \[\large \left| x \right|^{2}+\left| x+2 \right|^{2}=52\]

OpenStudy (anonymous):

is that the equation to figure out the problem

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok let try to work it out

OpenStudy (anonymous):

is it four

OpenStudy (anonymous):

yes, i get the same as wel

OpenStudy (anonymous):

cool ok thanks here is another problem Translate this problem to an equation and solve. The area of a rug is 36 square feet. The length of the rug is 5 feet longer than the width. What is the width of the rug? feet?

OpenStudy (anonymous):

Alright let me think

OpenStudy (anonymous):

ok will do and ill be trying too

OpenStudy (anonymous):

I guess you'll need to use the fact that \[Area = W*L\] and L = W+5

OpenStudy (anonymous):

the area is length*breadth length=5+breadth so area =breadth*(breadth+5) 36=b^2+5b so solving it we get b=4 this is wat i got

OpenStudy (anonymous):

Let the even numbers be x and x+2 \[x ^{2}+\left( x+2 \right)^{2}=52, where x>0\] \[x ^{2}+x ^{2}+4x+4=52\] \[2x ^{2}+4x-48=0\] \[x ^{2}+2x-24=0\] \[x ^{2}+6x-4x-24=0\] \[x \left( x+6 \right)-4\left( x+6 \right)=0\] \[\left( x+6 \right)\left( x-4 \right)=0\] x+6=0,x=-6 (rejected,xx>0) or x-4=0 x=4 smallest numbers= 4

OpenStudy (anonymous):

seems right

OpenStudy (anonymous):

Translate this problem to an equation and solve. The area of a square garden is 81 square feet. What is the length of each side?

OpenStudy (anonymous):

\[\sqrt{81}\]

OpenStudy (anonymous):

oh you need to translate it to an equation first, let's see.

OpenStudy (anonymous):

x*x=81 so x²=81

OpenStudy (anonymous):

its 9 cuase it is a square garden so that makes it a 9x9 garden right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Translate this problem to an equation and solve. The area of a patio is 72 square feet. The length of the patio is 1 foot longer than the width. What is the width of the patio?

OpenStudy (anonymous):

i square garden by definition has an area that is the square of the sides, so the sides are the root of the area.

OpenStudy (anonymous):

Same s the rug problem

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8 is the width right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The area of a rug is 66 square feet. The length of the rug is 5 feet longer than the width. What is the width of the rug? feet?

OpenStudy (anonymous):

simular question as the previous